# A Large-Scale Privacy Assessment of Android Third-Party SDKs

### v1

@@ 8.3. Limitations
While our static taint analysis approach is effective, dynamic testing can improve efficiency by automatically identifying and excluding logically invalid tainted traces, reducing manual validation efforts. However, dynamic SDK testing is challenging due to inadequate documentation, undisclosed interfaces, and highly event-driven/credential-required logic. We aim to explore potential dynamic testing solutions in the future.
@@
Scope of SDKs
Our assessment covered 158 Android SDKs from two mainstream release platforms. However, privacy data leakage can occur through non-Android APIs and even pure Java logic, suggesting our future work should extend beyond Android SDKs to any libraries or tools integrated into Android apps capable of handling personal data.
9. Related Works
@@
Numerous works focus on detecting inconsistencies between the privacy practices of an application and its descriptions or its privacy policies (bui2023detection, ; gorla2014chabada, ; liu2018large, ; pandita2013whyper, ; qu2014autocog, ; slavin2016toward, ; wang2018guileak, ; zhou2022policycomp, ). To automate the risk assessment of Android applications, WHYPER (pandita2013whyper, ) and AutoCog (qu2014autocog, ) leverage NLP techniques to evaluate the necessity of required permissions based on applications’ descriptions. Gorla et al. (gorla2014chabada, ) proposes a prototype to identify mismatches between app descriptions and app behaviors by clustering Android apps based on their description topics and identifying outliers in each cluster with respect to their API usage.
More recently, there have been several works examining third-party data harvesting, user consent for data sharing, service misconfigurations, and privacy policy conflicts in Android apps. XFinder (wang2021understanding, ) uses dynamic analysis and NLP techniques to parse terms-of-service and extract data-sharing policies. Nguyen et al. (nguyen2021share, ) studies user consent verification through traffic analysis and ablation experiments. Zhang et al. (zhang2020does, ) highlights the issue of analytical service misconfigurations, while Yu et al. (yu2018ppchecker, ) scrutinizes the apps’ privacy policies with regard to the inclusion of third-party libraries. Andow et al. (andow2020actions, ) proposes POLICHECK that identifies ambiguous or omitted disclosures of third parties. Despite these efforts, existing tools often neglect to assess the compliance of third-party libraries themselves with regulatory requirements. In comparison, our work focuses on compliance of third-party SDKs, checking the practical end-to-end data handling against declarations of developers. We also specifically consider the elevated risk of privacy leakage in the context of pre-installed apps.
@@
There is a substantial body of literature addressing various aspects of privacy concerns in the Android ecosystem (enck2014taintdroid, ; zhang2009peep, ; zhou2013identity, ). Recently, Meng et al. (meng2023post, ) studied user-unresettable identifier safeguards on a wide range of Android devices. They unveil widespread privacy issues, emphasizing the necessity for robust device-level data protection regulations. He et al. (he2018dynamic, ) utilizes dynamic analysis to explore privacy leaks from third-party libraries in Android apps, focusing on permission-related personal data. Their methods achieve a high real-time detection accuracy across 150 popular apps. Ekambaranathan et al. (ekambaranathan2021money, ) focuses on children’s app data usage and disclosure, gaining insight from surveys and interviews with app developers. Liu et al. (liu2019privacy, ) examines data leakage from nine analytic libraries across 300 apps, using both static and dynamic analyses. Razaghpanah et al. (razaghpanah2018apps, ) detects third-party advertising and tracking services via dynamic analysis of network traffic data. However, it identifies limited types of personal data and hardware identifiers. Our work comprehensively investigates privacy compliance at the SDK level on a large scale, complementing existing research to unveil the privacy preservation landscape in the software supply chain of the Android ecosystem.
10. Conclusion
@@
In this study, we have delved deep into the intricate privacy challenges associated with the widespread use of third-party SDKs in Android apps. Our large-scale analysis involved an evaluation of 158 third-party SDKs, prevalent across Android applications, with an approach intertwining rigorous taint analysis for data collection practices and LLM-aided interpretation of privacy policies. The findings were startling, with 338 potential privacy leakages detected from 158 SDKs. Besides that, our study reveals that less than 70% of examined SDKs provide privacy policies, among which approximately 37% of SDKs are found indulging in the over-collection of privacy data, signaling a clear violation of privacy norms. Based on our findings, we believe that tighter regulatory measures, improved development guidelines, and transparent permission management are essential steps toward mitigating privacy risks in Android applications.
Appendix A Appendix
A.1. The Five Categories of Privacy Data
(C1) Chip, Cellular & Peripheral
Each Android smartphone device has several unique identifiers that are associated with its hardware components including the motherboard, cellular chip, and peripheral hardware (e.g., external storage and SD card). These identifiers, such as serial, IMEI, and IMSI, are usually user-unresettable. Moreover, the invocation of some core functionalities associated with the cellular chip are also considered privacy sensitive (e.g., dialing a phone call, and sending an SMS), and therefore are categorized in this class, too. Apps’ access to these data often requests privileged permissions. Starting from Android 10, all user-unresettable device identifiers need “READ_PRIVILEGED_PHONE_STATE” permission, and therefore, are supposed to be read by system apps only (google2023privacy10, ).
(C2) Wireless Communication
. WiFi and Bluetooth are two common features of Android smartphones to perform wireless communication. Privacy data of this class include persistent identifiers such as MAC addresses, environmental volatile information that can be used to infer the user’s identity, including paired Bluetooth device ID (shown as Bluetooth ID), IP address, SSID and BSSID, and privacy-sensitive data generated via Wifi or Bluetooth functionalities, for example, making a call through Internet (SIP) or Bluetooth. In addition, since Android started third-party apps’ support of ultra-wideband (UWB) communication in version 12 and above (google2023uwb, ), we also consider data generated by UWB in this class. Considering data in this category may partially disclose users’ identity and/or location, Android has strengthened the permission mechanism for API requesting this type of data. For example, the MAC address randomization has been introduced in Android 6 and escalated restriction has been further enforced in Android 10 (google2023privacy10, ).
(C3) Camera, Location & Other Sensors
A modern Android smartphone is usually equipped with many sensors, for example, the camera, location sensor (e.g., GPS), and motion sensor. The data produced by these sensors are widely recognized as user privacy because they can be used to infer a user’s behavior, customs, and environment.
(C4) Media & Software Specific
This class concerns data associated with multimedia features and user ID generated by the OS or certain software. For example, the root directory of media files can be misused by malicious apps to save and read irrelevant files, audio recording, and screenshot functions can leak the user’s biometric voice and behavior. The software-defined ID can uniquely identify the user himself/herself but it can be reset by users. It is noteworthy that the list of apps or packages installed has been identified as privacy sensitive and can only be accessed by system apps since Android 11 (google2023privacy11, ).
(C5) Personal Data
Data falling in this class are mainly privacy-sensitive data that disclose users’ behavior or social relationships. For example, leakages of account information may disclose the social apps that the user has been logged in to. Unauthorized access to the contact list leaks users’ social relationships and can be manipulated for criminal purposes. Disclosure of clipboard data partially unveils the user’s real-time behavior, and therefore, severely infringes his/her privacy. Unlike other categories, permissions to access data in this group are highly varied. Accessing each type of data in this category usually requires unique permission.
A.2. List of Hypotheses Composed for NLI Tasks
Table 8. Combination of objects and applicable verbal phrases to form hypotheses
The complete list can be found in Table 8. Note that we compose a positive hypothesis (“It will…”), a contradiction hypothesis (“It does not…”), and an irrelevant hypothesis (“It does not mention whether to…”) for each combination of object and applicable verbal phrase. Given 169 combinations produced, we compose a total number of 507 hypotheses for the NLI tasks.
A.3. Discussion of False Positive Cases
In this section, we elaborate more on false positive (FP) cases observed through our privacy policy analysis. We take the three data types with the highest number of false positives as examples and discuss the potential reasons for the false positive results.
Account info
The account information in this study refers to certain sensitive personal data that constitutes an Android user’s account, such as the username and email address when the user sets up his/her mobile device for the first time. According to our analysis outcomes, account information ranks at the top by the number of false positive cases observed.
One possible reason for these FPs may be the too-broad scope of account information based on the knowledge of pre-trained LLMs. For example, many third-party apps or services, e.g., social networks and banking apps, request users to create their own accounts. SDKs connecting those apps and services may access domain-specific account details rather than the Android user information, which our adopted LLM usually fails to distinguish these two types of data. One statement in the privacy policy of the SDK “clevertap.android.sdk” could support this explanation, as it describes “personal info” to be collected as “The way users refer to themselves, such as their first or last name, or nickname” in its privacy policy. Our NLI model-based analysis determines that the Android account information is mentioned although the “first or last name” and “nickname” only refer to the user details of the in-house services.
Another reason leading to the high FP occurrences might be because the pre-trained LLMs tend to hold a preservative stance when considering account information and adopt a too broad sense of personal data into this scope. For example in our analysis of the privacy policy of SDK “com.razorpay.checkout”, our adopted model produces a high confidence score when determining whether the sentence “Device or other IDs, Device or other IDs” implies users’ account information. In that case, our adopted model likely considers the account information under the scope of “other IDs” mentioned in the policy text.
Network info
The network info in this study refers to the diverse data of the connected network configuration, which could be jointly used for user profiling purposes. However, this term may be too coarse and ambiguous for the LLMs. FPs in this data type are caused by our adopted model’s misunderstanding, which mistakenly considers any explicit usage of the network or Internet as an access to network information. For example, in an SDK named “com.zendesk.chat”, a statement about collecting users’ location “Location is not resolved on the device, but it is resolved on the backend based on the client’s IP address” is detected to collect network information, too.
SMS
The SMS in this paper refers to the data obtained through invoking SMS-related resources and functionalities, contributing to the third-highest number of FP cases. Our manual analysis finds that the adopted LLM manages to understand the exact literal meaning of SMS. However, we suspect that our model behaves too preservative to extend the scope of SMS to all sorts of messages. For example, during our analysis of the SDK “com.zendesk.messaging”, we find that the LLM misunderstands the difference between e-mail and SMS messages. It produces a positive prediction for collecting SMS although the privacy policy means to send e-mails to the users for customer service purposes.
A.4. Writing into System Settings and Pre-installed Apps
Writing into System Settings
According to the Android developer documentation, an app is requested to pose WRITE_SETTINGS permission to add or modify an entry into the system settings. That permission is under the protection level signature. For that reason, only apps signed with the platform signature can obtain that protection level and proceed to system settings modification. These apps also have to be system apps, e.g., installed in “/system/app/” or “/system/priv-app/” directory, to request the WRITE_SETTINGS permission.
However, our manual inspection of AOSP source code finds that the aforementioned permission is not the only way to enable an app to write or modify system settings. When an app attempts to write into system settings, the OS will check the protection level that the requesting app has attained. If any one of the following five tags is found from the app’s ApplicationInfo metadata, namely “signature” (with the tag value of 2), “appop” (40), “pre23” (80), “preinstalled” (400), and “role” (4000000), the OS will allow the requesting app to write into the system settings. Among the five flags, the preinstalled flag is set to all pre-installed apps by the OS.
Pre-installed Apps
According to Google’s documentation for system developers, there is no restriction posed to system developers in selecting pre-installed apps on a particular device. In other words, pre-installed apps do not have to apps with shared uid (i.e., user ID) and/or signed with the platform signature, such as android.uid.shell, android.uid.bluetooth, android.uid.phone, etc., but also be arbitrary third-party apps despite this practice is not encouraged by Google (google2024system, ).
According to the definition of protection level in AOSP source code (you may refer to the code “R.attr#ProtectionLevel” in the AOSP repository), “preinstalled” is a flag that is automatically granted to an arbitrary app pre-installed on the system image. An app with a “preinstalled” flag is given privilege to perform certain sensitive operations. Apart from writing into the system settings, there are some other privileged permissions can be automatically granted to pre-installed apps, including but not limited to AMBIENT_WALLPAPER, MANAGE_EXTERNAL_STORAGE, MANAGE_MEDIA,
READ_DEVICE_CONFIG, etc.
Note that pre-installed apps are not equivalent to “preloaded apps”. Preloaded apps can be installed in an ordinary location (e.g., /data/app/) or system directory (e.g., ‘‘/system/app/’’ and ‘‘/system/priv-app/’’), but pre-installed apps must be installed in a system directory. Generally, pre-installed apps stand for preloaded apps granted “preinstalled” privilege.
A.5. List of Tested Devices
No. Manufacturer Model
1 Xiaomi Xiaomi 11
2 Oppo Find X3
3 Vivo X70 Pro +
4 Honor Honor 50
5 OnePlus OnePlus 9
6 Samsung S21
7 Huawei P40
8 Lenovo Z6
9 Google Pixel Pixel 4 (US version)

### v2

@@ VI-C Limitations
While our static taint analysis approach is effective, dynamic testing can improve efficiency by automatically identifying and excluding logically invalid tainted traces, reducing manual validation efforts. However, dynamic SDK testing is challenging due to inadequate documentation, undisclosed interfaces, and highly event-driven/credential-required logic. For example, some finance and customer service SDKs require a valid license or token to function. Satisfying all functional prerequisites of the collected SDKs can be another challenging task and therefore is out of the scope of this study. We plan to explore potential dynamic testing solutions in the future.
@@
VII Ethics Considerations
We have reported all our findings to the relevant parties and we also have kept them confidential for at least 90 days. For each SDK that found a compliance issue, we manually searched the contact left by its developers and e-mailed our findings since our first round of SDK investigation. Although we observed slight mitigation from re-inspection, we did not receive positive responses regarding non-compliance concerns from the developers that we have attempted to contact. We also expressed our concern to Google about the system settings injection issues (to be detailed in Section V) and suggested removing the pre-installed privileges in writing system settings in the future release. Google replied and suggested that it is at the device manufacturer’s discretion to select and grant pre-installed privileges to third-party apps.
VIII Related Works
@@
Numerous works focus on detecting inconsistencies between the privacy practices of an application and its descriptions or its privacy policies [53, 54, 55, 56, 57, 58, 59, 60]. This line of research begins with automatic risk assessment of Android apps, such as WHYPER [56] and AutoCog [57] that leverage NLP techniques to evaluate the necessity of required permissions based on apps’ descriptions. Gorla et al. [54] study mismatches between app descriptions and app behaviors by clustering Android apps based on their description topics and identifying outliers in each cluster with respect to their API usage. Recent work extends the scope of privacy compliance evaluation to examining third-party data harvesting, user consent, service misconfigurations, and quality of privacy policies at the app level. XFinder [27] uses dynamic analysis and NLP techniques to parse terms-of-service and extract data-sharing policies. Nguyen et al. [61] study user consent verification through traffic analysis and ablation experiments. Zhang et al. [62] highlight the issue of analytical service misconfigurations, while Yu et al. [63] scrutinize the apps’ privacy policies with regard to the inclusion of third-party libraries. Andow et al. [32] propose POLICHECK that identifies ambiguous or omitted disclosures of third parties. Pan et al. [64] study the automated generation of privacy policies for mobile apps and unveil that defective compliance widely exists in the automatically generated privacy policies. Zhao et al. [65] study privacy compliance of third-party libraries in Android through the lens of pattern matching in privacy disclosure. Compared with it, our work investigates SDKs, which are more complex in implementation and business logic. Our study also resorts to a language model in privacy disclosure analysis that does not request manual data annotation and model training. Parallel research by Xiao et al. [66] investigate compliance of third-party libraries by assessing the consistency between API documents and dynamic analysis results. In comparison, our work focuses on compliance of SDKs that contains more complicated implementation that hinders from effective dynamic analysis, and meanwhile our study examine the privacy disclosure by directly interpreting the privacy policies, which are written in natural language without a standard format.
@@
There is a substantial body of literature addressing various aspects of privacy concerns in the Android ecosystem [28, 67, 68]. Meng et al. [43] study user-unresettable identifier safeguards on a wide range of Android devices. He et al. [69] utilize dynamic analysis to explore the leakage of permission-related data from third-party libraries in Android apps Ekambaranathan et al. [70] investigate data usage and disclosure in children’s app. Liu et al. [71] examine data leakage from nine analytic libraries across 300 apps, using both static and dynamic analyses. Razaghpanah et al. [72] detect third-party advertising and tracking services via dynamic analysis of network traffic data.
IX Conclusion
@@
In this study, we investigate the intricate privacy challenges associated with the widespread use of third-party SDKs in Android apps. Our large-scale analysis involved an evaluation of 158 third-party SDKs, prevalent across Android applications, with an approach intertwining rigorous taint analysis for data collection practices and LLM-aided interpretation of privacy policies. The findings were startling, with 338 potential privacy leakages detected from 158 SDKs. Besides that, our study reveals that less than 70% of examined SDKs provide privacy policies, among which approximately 37% SDKs are found indulging in the over-collection of privacy data, signaling a clear violation of privacy norms. Our findings suggest the need for stricter regulations, improved development guidelines, and more transparent permission management to reduce privacy risks in Android apps.